table 11.1 Corporate Tax Rates Tax Rat Taxable Income $0 - $50,000 50,001 - 75,000 15% 75,001 100,000 100,001 - 335,000 335,0PDQ, Inc. expects EBIT to be approximately $11 million per year for the foreseeable future, and that it has 25,000 20-year, 8 percent annual coupon bonds outstanding. (Use Table 11.1.)

What would the appropriate tax rate be for use in the calculation of the debt component of PDQ’s WACC?
Answer #1

Taxable income = EBIT - interest expense.

Interest expense = bonds outstanding * face value * interest rate = 25,000 * 1000 * 8% = $2,000,000.

Taxable income = $11,000,000 - $2,000,000 = $9,000,000.

Tax expense = ($50,000 * 15%) + (($75,000 - $50,001) * 25%) + (($100,000 - $75,001) * 34%) + (($335,000 - $100,001) * 39%) + (($9,000,000 - $335,001) * 34%) = $3,059,999.

Average tax rate = tax expense / taxable income = $3,059,999 / $9,000,000 = 34.00%.

The appropriate tax rate is 34.00%.